OpenStudy (anonymous):

(a) Deandre runs 6 miles in 36 minutes. How many minutes does he take per mile? (b) At Cheng's Bike Rentals, it costs $14 to rent a bike for 4 hours. How many hours of bike use does a customer get per dollar? If necessary, round your answers to the nearest hundredth.

OpenStudy (calculusxy):

(A) you can divide 36 minutes per six miles. Resulting in 6 minutes per mile.

OpenStudy (anonymous):

B ?

OpenStudy (anonymous):

hours per dollar = hours/dollars = 4/14 = 2/7. use calculator and round :)

OpenStudy (anonymous):

Round 2/7 ??

OpenStudy (anonymous):

use calculator and divide 2 by 7. round your result.

OpenStudy (anonymous):

0.28

OpenStudy (anonymous):

so it will be 0.30?